# Locally homogeneous Axiom A flows II: geometric structures for Anosov subgroups

### v2

@@ 8.3.1. The original definition of Anosov subgroups
[…]
The equivalence of Definition 7.1 with [GW12, Def. 2.10, Rem. 2.11], whose proof goes through [KP22] and [KLP18b, BPS19], gives the following property of $\Theta$-Anosov subgroups:
@@ 8.3.2. The tangent bundle of the flow space
Given a point $\mu=(\mu^{+},\mu^{-})\in{\mathbb{L}}$, denote by $([\mu^{+}],[\mu^{-}])\in{{\mathcal{F}}^{\pitchfork}}$ its projection, and by $\mathfrak{n}^{+}\subset\mathfrak{g}$ (resp. $\mathfrak{n}^{-}\subset\mathfrak{g}$) the nilpotent radical of the parabolic Lie algebra of the stabilizer of $[\mu^{+}]\in{\mathcal{F}}^{+}$ (resp. $[\mu^{-}]\in{\mathcal{F}}^{-}$). Using once again the notation $\overline{X}(\mu)=\left.\frac{d\,}{dt}\right|_{t=0}\exp(tX)\cdot\mu$ for the fundamental vector field of $X\in\mathfrak{g}$, we get a grading of the tangent space $T_{\mu}{\mathbb{L}}$ as
@@
Restricting this structure to ~$\widetilde{\mathcal{M}}\subset\mathbb{L}$ and taking the quotient by the $\Gamma$-action, the grading (8.4) descends yielding a flow invariant splitting
@@ 8.3.3. Proof of hyperbolicity
The flow invariance $\pi_{\pm}\circ\phi^{t}=\pi_{\pm}$, when differentiated, shows that the isomorphism (8.5) intertwines the flows. The $G$-equivariant vector bundles $\pi_{\pm}^{*}T{\mathcal{F}}^{\pm}\to{\mathbb{L}}$ can be restricted to $\Gamma$-invariant bundles over ~$\widetilde{\mathcal{M}}\subset{\mathbb{L}}$ and descended to vector bundles $\mathcal{T}^{\pm}$ over $\mathcal{M}$ each equipped with a 1-parameter group of vector bundle isomorphisms still denoted by $\varphi^{t}_{\pm}$. Upon restriction to $\mathcal{K}\subset\mathcal{M}$, we obtain a commutative diagram
@@
[…]
@@ 9.1.1. The universal para-Kähler structure
A para-Kähler structure on a smooth manifold $M$ of dimension $2k$ is a tuple $(\langle-,-\rangle,I,\omega)$ where $\langle$-,$\rangle$ - , - ⟩ is a neutral signature $(k,k)$ pseudo-Riemannian metric on $M$, the bundle homomorphism $I:TM\rightarrow TM$ is parallel for the Levi-Civita connection of $\langle$-,$\rangle$ - , - ⟩ and idempotent (i.e. $I^{2}=\mathrm{Id}$), and $\langle I(-),-\rangle=\omega$ is a symplectic form. It is the analog of a Kähler structure where the complex numbers are replaced by the paracomplex numbers $a+b\varepsilon$ where $a,b\in\mathbb{R}$ and $\varepsilon^{2}=1.$ The ±$\pm 1$-eigenspaces of $I:TM\rightarrow TM$ constitute a pair of transverse, integrable Lagrangian distributions for the symplectic form $\omega$, and the para-Kähler structure is equivalent to a symplectic manifold with a pair of transverse Lagrangian foliations.
@@
can be used to push forward the Killing form $\mathrm{B}_{\bf\mathfrak{g}}$ to a $G$-invariant pseudo-Riemannian metric $\langle-,-\rangle_{{{\mathcal{F}}^{\pitchfork}}}$ of signature $(n,n)$ on ${\mathcal{F}}^{\pitchfork}$ (where $n=\dim{\mathcal{F}}^{+}=\dim{\mathcal{F}}^{-}$) [^4] This is the same structure as defined in Section 3.3.
@@
for which the subbundles ±$\mathcal{L}^{\pm}\subset T{{\mathcal{F}}^{\pitchfork}}$ are totally isotropic. Note that since the Lie subalgebras $\mathfrak{n}^{+},\mathfrak{n}^{-}$ satisfy $[\mathfrak{n}^{+},\mathfrak{n}^{-}]\subset\mathfrak{l}$, where $\mathfrak{l}$ denotes the Lie algebra of the Levi stabilizer of $\mathbf{x}$, we have that
@@
and find that $\omega_{{\mathcal{F}}^{\pitchfork}}$ is skew-symmetric and non-degenerate (because the Killing form induces a non-degenerate pairing $\mathrm{B}_{\bf\mathfrak{g}}:\mathfrak{n}^{+}\times\mathfrak{n}^{-}\to\mathbb{R}$). Cartan’s magic formula applied to (9.6) shows that $d\omega_{{\mathcal{F}}^{\pitchfork}}=0$, i.e. that $\omega_{{\mathcal{F}}^{\pitchfork}}$ is a symplectic form, so we have defined a $G$-invariant para-Kähler manifold $({\mathcal{F}}^{\pitchfork},\langle-,-\rangle_{{\mathcal{F}}^{\pitchfork}},I_{{\mathcal{F}}^{\pitchfork}},\omega_{{\mathcal{F}}^{\pitchfork}})$.
@@ 9.1.3. Period functions and closed 2-forms
Consider a transverse flag manifold ${{\mathcal{F}}^{\pitchfork}}={\mathcal{F}}^{+}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}^{-}$, and recall from Definition 4.6 its space $\mathcal{P}({{\mathcal{F}}^{\pitchfork}})$ of period functions of ${{\mathcal{F}}^{\pitchfork}}$, i.e. real valued $G$-invariant functions on $\mathrm{Stab}_{{\mathcal{F}}^{\pitchfork}}=\left\{(\mathbf{x},g)\in{{\mathcal{F}}^{\pitchfork}}\times G\,\middle|\,g\cdot\mathbf{x}=\mathbf{x}\right\}$ that restrict to a group homomorphism on each fiber.
@@
The coadjoint orbit $\mathcal{O}_{\ell}$ is always hyperbolic: the Killing dual of the linear form $d\ell(\mathbf{x})$ at any $\mathbf{x}\in{{\mathcal{F}}^{\pitchfork}}$ must lie in $\mathfrak{z}_{\mathrm{split}}(\mathfrak{l})$, because $d\ell(\mathbf{x})$ vanishes on the Lie algebra of the maximal compact torus of $\mathfrak{l}$ as well as on the derived algebra $[\mathfrak{l},\mathfrak{l}]$. In particular, $\mathcal{O}_{\ell}$ is always $G$-equivariantly diffeomorphic to some transverse flag space.
@@
If $\ell\in\mathcal{P}({{\mathcal{F}}^{\pitchfork}})$ is regular, then $\langle-,-\rangle_{\ell}:=\omega^{\ell}(I_{{\mathcal{F}}^{\pitchfork}}(-),-)$ is a pseudo-Riemannian metric of neutral signature, defining a (generally) different para-Kähler structure on ${{\mathcal{F}}^{\pitchfork}}$.
@@ 9.2.2. The canonical 1-form
$\varphi_{\hat{\mathbf{x}}}:\left\{\begin{array}[]{ccc}G&\to&{\mathbb{L}}\\ g&\mapsto&g\cdot\hat{\mathbf{x}}.\end{array}\right.$
@@ 9.3.1. Regularity and multi-densities
Considering opposite flag manifolds ${\mathcal{F}}^{+},{\mathcal{F}}^{-}$ corresponding to subsets $\Theta,\iota_{o}(\Theta)\subset\Delta$ and $\mathbf{s}\in\mathbb{R}^{\Theta}$, it is particularly easy to tell when the flow space ${\mathbb{L}}_{\Theta}^{\mathbf{s}}$ over ${{\mathcal{F}}^{\pitchfork}}={\mathcal{F}}^{+}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}^{-}$ is regular, i.e. when its period function is regular.
@@ 9.3.2. Contact realization of refraction flows
Consider some transverse flag space ${{\mathcal{F}}^{\pitchfork}}$ and the associated subset $\Theta\subset\Delta$. Let $\Gamma<G$ be a torsion-free $\Theta$-Anosov subgroup, and $\mathbf{s}\in\mathbb{R}^{\Theta}$ a $(\Gamma,\Theta)$-admissible parameter. If $s_{\alpha}\neq 0$ for all $\alpha\in\Theta$, then the flow $\phi^{t}$ on the quotient manifold $\mathcal{M}_{\Gamma,\Theta}^{\mathbf{s}}$ is the Reeb flow of a contact form.
@@
${{\mathbb{L}}^{\mathbf{s}}_{\Theta}}$${{\mathbb{L}}_{\Theta^{\prime}}^{\mathbf{s}^{\prime}}}$${{{\mathcal{F}}^{\pitchfork}}}$${{\mathcal{F}}_{\Theta^{\prime}}^{\mathbf{s}^{\prime}}}$$\scriptstyle{p}$$\scriptstyle{\widetilde{f}}$$\scriptstyle{p^{\prime}}$$\scriptstyle{\pi}$
The transverse limit maps $\xi^{\pitchfork}:\partial_{\infty}\Gamma^{(2)}\to{{\mathcal{F}}^{\pitchfork}}$ and $\xi_{\Theta^{\prime}}^{\pitchfork}:\partial_{\infty}\Gamma^{(2)}\to{\mathcal{F}}_{\Theta^{\prime}}^{\pitchfork}$ satify $\xi_{\Theta^{\prime}}^{\pitchfork}=\pi\circ\xi^{\pitchfork}$ (because of Lemma 3.40). This implies that $\widetilde{f}$ maps ~$\widetilde{\mathcal{M}}^{\mathbf{s}}_{\Gamma,\Theta}$ onto ~$\widetilde{\mathcal{M}}^{\mathbf{s}^{\prime}}_{\Gamma,\Theta^{\prime}}$ and $\widetilde{\mathcal{K}}^{\mathbf{s}}_{\Gamma,\Theta}$ onto $\widetilde{\mathcal{K}}^{\mathbf{s}^{\prime}}_{\Gamma,\Theta^{\prime}}$, inducing a flow-equivariant surjective submersion $f:\mathcal{M}_{\Gamma,\Theta}^{\mathbf{s}}\to\mathcal{M}_{\Gamma,\Theta^{\prime}}^{\mathbf{s}^{\prime}}$ that maps $\mathcal{K}^{\mathbf{s}}_{\Gamma,\Theta}$ onto $\mathcal{K}^{\mathbf{s}^{\prime}}_{\Gamma,\Theta^{\prime}}$. This restriction is also injective because of the flow-equivariance, so it must be a homeomorphism by compactness of $\mathcal{K}^{\mathbf{s}}_{\Gamma,\Theta}$. ∎
@@ 10.1. The density multiflow space
Consider opposite flag manifolds ${\mathcal{F}}^{+},{\mathcal{F}}^{-}$ associated to subsets $\Theta,\iota_{o}(\Theta)\subset\Delta$ respectively, and let ${{\mathcal{F}}^{\pitchfork}}={\mathcal{F}}^{+}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}^{-}$ be the transverse flag space. Instead of considering multidensities on ${\mathcal{F}}^{+}$ and ${\mathcal{F}}^{-}$ as we did for the flow spaces in Section 6, we will keep track of individual densities on each simple flag manifold dominated by ${\mathcal{F}}^{+}$ and ${\mathcal{F}}^{-}$. Recall from Corollary 3.20 that for every $\alpha\in\Theta$ we have a $G$-equivariant projection $\pi_{\alpha}:{{\mathcal{F}}^{\pitchfork}}\to{\mathcal{F}}^{\pitchfork}_{\alpha}={\mathcal{F}}_{\alpha}\,{\scriptscriptstyle\stackrel{{\scriptstyle\scriptscriptstyle\pitchfork}}{{\scriptstyle\times}}}\,{\mathcal{F}}_{\iota_{o}(\alpha)}$.
@@
The multiflow space over ${{\mathcal{F}}^{\pitchfork}}$ is the $\mathbb{R}^{\Theta}$-principal bundle
@@
${\mathbb{W}_{\Theta}}$${{\mathbb{L}}_{\Theta}^{\mathbf{s}}}$${{{\mathcal{F}}^{\pitchfork}}}$$\scriptstyle{\Pi_{\Theta}^{\mathbf{s}}}$$\scriptstyle{\hat{p}}$$\scriptstyle{p}$
@@
$=\phi^{\mathbf{t}\cdot\mathbf{s}}\left(\Pi_{\Theta}^{\mathbf{s}}(\nu)\right)\,.\vspace*{-1em}$
@@
Consider some $(\Gamma,\Theta)$-admissible parameter $\mathbf{s}$ (e.g. $s_{\alpha}=1$ for every $\alpha$). Then $\Pi_{\Theta}^{\mathbf{s}}$ maps $\hat{p}^{-1}(\Omega_{\Gamma})$ onto ~$\widetilde{\mathcal{M}}^{\mathbf{s}}_{\Gamma,\Theta}$ in a $\Gamma$-equivariant way, so the proper discontinuity on ~$\widetilde{\mathcal{M}}^{\mathbf{s}}_{\Gamma,\Theta}$ implies the same property on $\hat{p}^{-1}(\Omega_{\Gamma})$. ∎
@@
We can therefore consider the quotient manifold $\mathcal{N}_{\Gamma,\Theta}:=\Gamma\backslash\hat{p}^{-1}(\Omega_{\Gamma})$. The equivariance formula in Lemma 10.2 shows that $\mathcal{N}_{\Gamma,\Theta}$ (resp. $\widehat{\mathcal{K}}_{\Gamma,\Theta}:=\Gamma\backslash\hat{p}^{-1}(\Lambda^{\pitchfork}_{\Gamma})$) is a principal $\mathbf{s}^{\perp}\simeq\mathbb{R}^{\lvert\Theta\rvert-1}$-bundle over $\mathcal{M}^{\mathbf{s}}_{\Gamma,\Theta}$ (resp. $\mathcal{K}^{\mathbf{s}}_{\Gamma,\Theta}$). In particular, $\widehat{\mathcal{K}}_{\Gamma,\Theta}$ is never compact when $\lvert\Theta\rvert\geq 2$.
@@ 11.2. Lagrangians in $(E,\omega)$
$\Lambda_{\Gamma}^{\pitchfork}$ =$=\left\{(\xi(z^{+}),\xi(z^{-}))\,\middle|\,z^{+}\neq z^{-}\right\}\subset\mathrm{Lag}^{\pitchfork}\,,$
$\Omega_{\Gamma}$ =$=\left\{(L_{1},L_{2})\in\mathrm{Lag}^{\pitchfork}\,\middle|\,\forall z\in\partial_{\infty}\Gamma~{}L_{1}\pitchfork\xi(z)\textrm{ or }\xi(z)\pitchfork L_{2}\right\}$
lift respectively to invariant subsets $\widetilde{\mathcal{K}}_{\Gamma}\subset\widetilde{\mathcal{M}}_{\Gamma}\subset{\mathbb{L}}$ where the action of $\Gamma$ on ~$\widetilde{\mathcal{M}}_{\Gamma}$ is properly discontinuous and the restriction to $\widetilde{\mathcal{K}}_{\Gamma}$ is cocompact, leading to a real analytic contact Axiom A flow on the quotient $\mathcal{M}_{\Gamma}=\Gamma\backslash\widetilde{\mathcal{M}}_{\Gamma}$ with non-wandering set $\mathcal{K}_{\Gamma}=\Gamma\backslash\widetilde{\mathcal{K}}_{\Gamma}$.
@@
It is possible to use non-zero volume forms $\omega_{i}\in\Lambda^{d}L_{i}^{*}$ instead of densities, but in the absence of a natural orientation of a Lagrangrian subspace, one has to consider them up to multiplication by $-1$. This leads either to this density construction, or to considering the Plücker embedding $\mathrm{Lag}\subset\mathcal{G}_{d}(E)\hookrightarrow\mathbb{P}(\Lambda^{d}E)$ then using the flow considered in [DMS24].

### v3

@@ 8.3.1. The original definition of Anosov subgroups
$V^{+}$$V^{+}$$V^{-}$$V^{-}$$\mathfrak{R}_{\sigma}$$\mathfrak{R}_{\sigma}$$\mathfrak{R}_{\sigma}$$\mathfrak{R}_{\sigma}$$\scriptstyle{\psi_{+}^{t}}$$\scriptstyle{\psi_{-}^{t}}$$\scriptstyle{\psi_{\sigma}^{t}}$$\scriptstyle{\psi_{\sigma}^{t}}$
The equivalence of Definition 7.1 with [GW12, Def. 2.10, Rem. 2.11], i.e. [KLP17, Theorem 1.1], gives the following property of $\Theta$-Anosov subgroups:
@@ 8.3.2. The tangent bundle of the flow space
Given a point $\mu=(\mu^{+},\mu^{-})\in{\mathbb{L}}$, denote by $([\mu^{+}],[\mu^{-}])\in{{\mathcal{F}}^{\pitchfork}}$ its projection, and by $\mathfrak{n}^{+}\subset\mathfrak{g}$ (resp. $\mathfrak{n}^{-}\subset\mathfrak{g}$) the nilpotent radical of the parabolic Lie algebra of the stabilizer of $\mu^{+}]\in{\mathcal{F}}^{+}$ (resp. $\mu^{-}]\in{\mathcal{F}}^{-}$). Using once again the notation $\overline{X}(\mu)=\left.\frac{d\,}{dt}\right|_{t=0}\exp(tX)\cdot\mu$ for the fundamental vector field of $X\in\mathfrak{g}$, we get a grading of the tangent space $T_{\mu}{\mathbb{L}}$ as
@@
Restricting this structure to $\widetilde{\mathcal{M}}\subset\mathbb{L}$ and taking the quotient by the $\Gamma$-action, the grading (8.4) descends yielding a flow invariant splitting
@@ 8.3.3. Proof of hyperbolicity
The flow invariance $\pi_{\pm}\circ\phi^{t}=\pi_{\pm}$, when differentiated, shows that the isomorphism (8.5) intertwines the flows. The $G$-equivariant vector bundles $\pi_{\pm}^{*}T{\mathcal{F}}^{\pm}\to{\mathbb{L}}$ can be restricted to $\Gamma$-invariant bundles over $\widetilde{\mathcal{M}}\subset{\mathbb{L}}$ and descended to vector bundles $\mathcal{T}^{\pm}$ over $\mathcal{M}$ each equipped with a 1-parameter group of vector bundle isomorphisms still denoted by $\varphi^{t}_{\pm}$. Upon restriction to $\mathcal{K}\subset\mathcal{M}$, we obtain a commutative diagram
@@
$V^{\pm}$$\mathcal{T}^{\pm}|_{\mathcal{K}}$$E^{\pm}|_{\mathcal{K}}$$\mathfrak{R}_{\sigma}$$\mathcal{K}$$\mathcal{K}$$\scriptstyle{\Phi_{\sigma}}$$\scriptstyle{\mathrm{Id}}$
@@ 9.1.1. The universal para-Kähler structure
A para-Kähler structure on a smooth manifold $M$ of dimension $2k$ is a tuple $(\langle-,-\rangle,I,\omega)$ where $\langle-,$-$\rangle$ is a neutral signature $(k,k)$ pseudo-Riemannian metric on $M$, the bundle homomorphism $I:TM\rightarrow TM$ is parallel for the Levi-Civita connection of $\langle-,$-$\rangle$ and idempotent (i.e. $I^{2}=\mathrm{Id}$), and $\langle I(-),-\rangle=\omega$ is a symplectic form. It is the analog of a Kähler structure where the complex numbers are replaced by the paracomplex numbers $a+b\varepsilon$ where $a,b\in\mathbb{R}$ and $\varepsilon^{2}=1.$ The $\pm 1$-eigenspaces of $I:TM\rightarrow TM$ constitute a pair of transverse, integrable Lagrangian distributions for the symplectic form $\omega$, and the para-Kähler structure is equivalent to a symplectic manifold with a pair of transverse Lagrangian foliations.
@@
[…]
@@
for which the subbundles $\mathcal{L}^{\pm}\subset T{{\mathcal{F}}^{\pitchfork}}$ are totally isotropic. Note that since the Lie subalgebras $\mathfrak{n}^{+},\mathfrak{n}^{-}$ satisfy $\mathfrak{n}^{+},\mathfrak{n}^{-}]\subset\mathfrak{l}$, where $\mathfrak{l}$ denotes the Lie algebra of the Levi stabilizer of $\mathbf{x}$, we have that
@@
[…]
@@ 9.1.3. Period functions and closed 2-forms
[…]
@@
The coadjoint orbit $\mathcal{O}_{\ell}$ is always hyperbolic: the Killing dual of the linear form $d\ell(\mathbf{x})$ at any $\mathbf{x}\in{{\mathcal{F}}^{\pitchfork}}$ must lie in $\mathfrak{z}_{\mathrm{split}}(\mathfrak{l})$, because $d\ell(\mathbf{x})$ vanishes on the Lie algebra of the maximal compact torus of $\mathfrak{l}$ as well as on the derived algebra $\mathfrak{l},\mathfrak{l}]$. In particular, $\mathcal{O}_{\ell}$ is always $G$-equivariantly diffeomorphic to some transverse flag space.
@@
[…]
@@ 9.2.2. The canonical 1-form
[…]
@@ 9.3.1. Regularity and multi-densities
[…]
@@ 9.3.2. Contact realization of refraction flows
[…]
@@
[…]
The transverse limit maps $\xi^{\pitchfork}:\partial_{\infty}\Gamma^{(2)}\to{{\mathcal{F}}^{\pitchfork}}$ and $\xi_{\Theta^{\prime}}^{\pitchfork}:\partial_{\infty}\Gamma^{(2)}\to{\mathcal{F}}_{\Theta^{\prime}}^{\pitchfork}$ satify $\xi_{\Theta^{\prime}}^{\pitchfork}=\pi\circ\xi^{\pitchfork}$ (because of Lemma 3.39). This implies that $\widetilde{f}$ maps $\widetilde{\mathcal{M}}^{\mathbf{s}}_{\Gamma,\Theta}$ onto $\widetilde{\mathcal{M}}^{\mathbf{s}^{\prime}}_{\Gamma,\Theta^{\prime}}$ and $\widetilde{\mathcal{K}}^{\mathbf{s}}_{\Gamma,\Theta}$ onto $\widetilde{\mathcal{K}}^{\mathbf{s}^{\prime}}_{\Gamma,\Theta^{\prime}}$, inducing a flow-equivariant surjective submersion $f:\mathcal{M}_{\Gamma,\Theta}^{\mathbf{s}}\to\mathcal{M}_{\Gamma,\Theta^{\prime}}^{\mathbf{s}^{\prime}}$ that maps $\mathcal{K}^{\mathbf{s}}_{\Gamma,\Theta}$ onto $\mathcal{K}^{\mathbf{s}^{\prime}}_{\Gamma,\Theta^{\prime}}$. This restriction is also injective because of the flow-equivariance, so it must be a homeomorphism by compactness of $\mathcal{K}^{\mathbf{s}}_{\Gamma,\Theta}$. ∎
@@ 10.1. The density multiflow space
[…]
@@
[…]
@@
[…]
@@
$=\phi^{\mathbf{t}\cdot\mathbf{s}}\left(\Pi_{\Theta}^{\mathbf{s}}(\nu)\right)\,.\vskip-10.00002pt$
@@
Consider some $(\Gamma,\Theta)$-admissible parameter $\mathbf{s}$ (e.g. $s_{\alpha}=1$ for every $\alpha$). Then $\Pi_{\Theta}^{\mathbf{s}}$ maps $\hat{p}^{-1}(\Omega_{\Gamma})$ onto $\widetilde{\mathcal{M}}^{\mathbf{s}}_{\Gamma,\Theta}$ in a $\Gamma$-equivariant way, so the proper discontinuity on $\widetilde{\mathcal{M}}^{\mathbf{s}}_{\Gamma,\Theta}$ implies the same property on $\hat{p}^{-1}(\Omega_{\Gamma})$. ∎
@@
[…]
@@ 11.2. Lagrangians in $(E,\omega)$
$\Lambda_{\Gamma}^{\pitchfork}$ $=\left\{(\xi(z^{+}),\xi(z^{-}))\,\middle|\,z^{+}\neq z^{-}\right\}\subset\mathrm{Lag}^{\pitchfork}\,$,
$\Omega_{\Gamma}$ $=\left\{(L_{1},L_{2})\in\mathrm{Lag}^{\pitchfork}\,\middle|\,\forall z\in\partial_{\infty}\Gamma~L_{1}\pitchfork\xi(z)\textrm{ or }\xi(z)\pitchfork L_{2}\right\}$
lift respectively to invariant subsets $\widetilde{\mathcal{K}}_{\Gamma}\subset\widetilde{\mathcal{M}}_{\Gamma}\subset{\mathbb{L}}$ where the action of $\Gamma$ on $\widetilde{\mathcal{M}}_{\Gamma}$ is properly discontinuous and the restriction to $\widetilde{\mathcal{K}}_{\Gamma}$ is cocompact, leading to a real analytic contact Axiom A flow on the quotient $\mathcal{M}_{\Gamma}=\Gamma\backslash\widetilde{\mathcal{M}}_{\Gamma}$ with non-wandering set $\mathcal{K}_{\Gamma}=\Gamma\backslash\widetilde{\mathcal{K}}_{\Gamma}$.
@@
It is possible to use non-zero volume forms $\omega_{i}\in\Lambda^{d}L_{i}^{*}$ instead of densities, but in the absence of a natural orientation of a Lagrangrian subspace, one has to consider them up to multiplication by 1-1. This leads either to this density construction, or to considering the Plücker embedding $\mathrm{Lag}\subset\mathcal{G}_{d}(E)\hookrightarrow\mathbb{P}(\Lambda^{d}E)$ then using the flow considered in [DMS24].